OpenStudy (anonymous):

Factor completely a^2-49

OpenStudy (mathstudent55):

Use the difference of two squares factoring. x^2 - y^2 = (x + y)(x - y)

OpenStudy (anonymous):

so then is the answer a(a-7^2)

OpenStudy (mathstudent55):

No. Follow the pattern. You start with x^2 - y^2 and you end up with (x + y)(x - y) In your case you start with a^2 - 49 = a^2 - 7^2, so you end up with (a + 7)(a - 7)

OpenStudy (anonymous):

so the the (a=7)(a-7) is the answer ok I understand thanks.